OpenStudy (anonymous):

sin 2(theta) + cos 2(theta) gives what

OpenStudy (mathteacher1729):

Are they asking you to simplify using the double-angle formula?

OpenStudy (anonymous):

they are asking to complete the formula just complete the formula

OpenStudy (anonymous):

= 1 (if they are in square form)

OpenStudy (anonymous):

@waterineyes it is double ang

OpenStudy (anonymous):

\[\huge \sin^ \theta + \cos^2 \theta = 1\]

OpenStudy (mathteacher1729):

Oh wait, is it \[\sin^2(\theta)+\cos^2(\theta)\] or \[\sin(2\theta)+\cos(2\theta)\]

OpenStudy (anonymous):

second one @mathteacher1729

OpenStudy (mathteacher1729):

Gotcha. Then it's the double-angle formula. All you have to do is find the double-angle formula (I recommend this cheat sheet http://tutorial.math.lamar.edu/cheat_table.aspx ) and then replace each term (sin and cos) with the appropriate expression.

OpenStudy (anonymous):

\[= 2\sin \theta \cos \theta + \cos^2 \theta - \sin^2 \theta\]

OpenStudy (apoorvk):

\(\sin2A = 2 \sin A\cos A\) \(\cos2A = \cos^2A - \sin^2A\) so, \(sin2A + cos 2A = 2 \sin A\cos A + \cos^2A - \sin^2A\) or, \(sin2A + cos 2A = (2 \sin A\cos A + \cos^2A + \sin^2A) - 2 \sin^2A\) or, \(sin2A + cos 2A = (\cos A + \sin A)^2 - 2 \sin^2A\)

OpenStudy (apoorvk):

I can't seem to think how this can be simplified further. Or may be the first step itself is the most simplified form possible.

OpenStudy (anonymous):

Moreover, \[\large = (cosA + sinA - \sqrt{2}sinA)(cosA + sinA + \sqrt{2}\sin A)\]

OpenStudy (apoorvk):

yeah, in factorised format^

OpenStudy (anonymous):

\[2\tan \theta + (1-\tan ^{2}\theta) \div (1+\tan ^{2}\theta )\]

OpenStudy (anonymous):

this is what she wanted

OpenStudy (anonymous):

She???

OpenStudy (anonymous):

miss

OpenStudy (anonymous):

Ok I got it now...

OpenStudy (anonymous):

i will give u one ques which u cant do at all

OpenStudy (anonymous):

Use these identities: \[\huge \color{green}{\sin2 \theta = \frac{2\tan \theta}{1 + \tan^2 \theta}}\] \[\huge \color{green}{\cos2 \theta = \frac{1 - \tan^2 \theta}{1 + \tan^2 \theta}}\]

OpenStudy (anonymous):

Just take the LCM: \[\large = \frac{2 \tan \theta + (1 - \tan^2 \theta)}{1 + \tan^2 \theta}\] Got it??

OpenStudy (anonymous):

i got it before u man

OpenStudy (anonymous):

Oh that is so nice then..